Circular No. 7012

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1998 QP_{54}

On Sept. 13, Roy A. Tucker (Tucson, AZ) reported his discovery of a comet in the course of his CCD asteroid-astrometry program with a 0.36m f/11 Schmidt-Cassegrain at the Goodricke-Pigott Observatory. G. V. Williams, Minor Planet Center, then identified the comet with 1998 QP_{54}, which had been reported by E. Bowell (observer W. D. Ferris, measurer B. W. Koehn) as an apparently quite ordinary minor planet in the course of LONEOS, the Lowell Observatory Near-Earth Object Search (cf. MPC 32398; the observations themselves, on Aug. 27 and 28, were published on MPS 2052). Williams then also identified single-night observations (also on Aug. 28) by LINEAR as belonging to the same object. Information about the object was then placed in The NEO Confirmation Page. In response, J. Tichá reported observations made at the Kleť Observatory that showed a faint coma (diameter 17′′) and a 35′′ tail in p.a. 226^{o}, and L. Šarounová reported from the Ondřejov Observatory on a narrow tail 5′ long in p.a. 135^{o} and only a small coma. Observations were also reported by T. Spahr from the Catalina Sky Survey. The LONEOS and Tucker discovery observations are listed here for reference, and the complete set of currently available observations is given on MPEC 1998-R27, which is also the source for the orbital elements and ephemeris below. The comet is clearly of short period, and it made a close approach to Jupiter early in 1992.

1998 UT α_{2000} δ_{2000} m_{2} Observer
Aug. 27.41419 0^{h}31^{m}51.^{s}89 +10^{o}28′55′′.0 17.1 Ferris
Sept. 13.21368 0 22 19.28 +15 12 15.4 16.7 Tucker
T = 1998 Oct. 6.698 TT  ω = 30.^{o}177
e = 0.55537  Ω = 342.001 }2000.0
q = 1.88548 AU  i = 17.825
a = 4.24060 AU n^{o} = 0.112866 P = 8.73 years
1998TT α_{2000} δ_{2000} Δ r ϵ β m_{2}
Sept. 14 0^{h}21.^{m}67 +15^{o}24.^{′}7 0.931 1.897 156.^{o}7 12.^{o}1 17.0
19 0 17.25 +16 40.2 0.917 1.893 160.0 10.4 16.9
24 0 12.42 +17 49.9 0.909 1.889 162.3  9.3 16.8
29 0 07.41 +18 52.9 0.906 1.887 163.1  8.9 16.8
Oct. 4 0 02.45 +19 48.6 0.909 1.886 162.0  9.4 16.8
9 23 57.76 +20 37.0 0.917 1.886 159.6 10.7 16.9
14 23 53.56 +21 18.3 0.931 1.887 156.2 12.3 17.0
1998 September 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden